### v1

# CoAR: Concept Injection into Autoregressive Models for Personalized Text-to-Image Generation

Abstract
The unified autoregressive (AR) model excels at multimodal understanding and generation, but its potential for customized image generation remains underexplored. Existing customized generation methods rely on full fine-tuning or adapters, making them costly and prone to overfitting or catastrophic forgetting. In this paper, we propose CoAR, a novel framework for injecting subject concepts into the unified AR models while keeping all pre-trained parameters completely frozen. CoAR learns effective, specific subject representations with only a minimal number of parameters using a Layerwise Multimodal Context Learning strategy. To address overfitting and language drift, we further introduce regularization that preserves the pre-trained distribution and anchors context tokens to improve subject fidelity and re-contextualization. Additionally, CoAR supports training-free subject customization in a user-provided style. Experiments demonstrate that CoAR achieves superior performance on both subject-driven personalization and style personalization, while delivering significant gains in computational and memory efficiency. Notably, CoAR tunes less than 0.05% of the parameters while achieving competitive performance compared to recent Proxy-Tuning. Code: https://github.com/KZF-kzf/CoAR.
Figure 1: Visualization of samples generated by our proposed CoAR, showcasing subject personalization (left), where a few reference images are provided, and style personalization (right), where the subject is rendered in user-specified artistic styles without additional training.
Introduction
Subject-driven image generation (Chen et al. 2023b; Ruiz et al. 2023; Li, Li, and Hoi 2023; Shin et al. 2025) aims to generate images of a specific subject or concept in new scenes, given only a few reference images. This capability has found applications in virtual try-on (Ye et al. 2025; Zhang et al. 2024), personalized content creation (Liu et al. 2024b), and advertising design (Chen et al. 2023b).
Most of the existing approaches, such as DreamBooth (Ruiz et al. 2023) and its extensions (Chen et al. 2023a; Han et al. 2023; Kumari et al. 2023), customize generation by adapting the distribution of pre-trained diffusion models to match the subject distribution through full fine-tuning or parameter-efficient techniques like LoRA (Hu et al. 2022). These methods typically associate the subject with a unique identifier [V], enabling personalized generation via prompt injection during inference. However, such techniques are predominantly designed for diffusion models (Esser et al. 2024; Rombach et al. 2022; Podell et al. 2023) and are not directly transferable to multimodal autoregressive (AR) models. Applying these methods naively to AR models poses two challenges.
  • Risk of overfitting. AR models are pretrained on large-scale multimodal corpora and expected to support both understanding and generation. In contrast, the target subject typically has only a few reference images (3–5). Such scarce data makes it extremely difficult to adapt the pre-trained models to the subject distribution, resulting in a risk of overfitting and catastrophic forgetting of the pre-trained knowledge.
  • Inefficiency. Extensively adjusting the pretrained model parameters to learn a single concept with the limited number of images per subject is intuitively inefficient.
We argue that instead of fine-tuning the pretrained model on limited data to shift its distribution toward a target subject, it is more effective to inject subject concepts directly into the model’s existing distribution, thereby fully leveraging the generative capabilities of large models. To this end, we introduce Concept Injection for AR Models (CoAR), an extremely efficient framework for customized generation in multimodal AR models. CoAR adopts a Layer-wise Multimodal Context Learning (LMCL) strategy, encoding subject-specific information through a small number of learnable prompts without modifying any pretrained parameters. To guide learning while preserving the model’s original capabilities, we propose a Dual Prior Preservation (DPP) loss that constrains training. Additionally, we design a Context-Aware Self-Regularization (CASR) loss to improve re-contextualization stability and output quality. Built upon this foundation, CoAR also supports training-free customization, enabling arbitrary subject–style combinations during inference while maintaining strong re-contextualization performance. With a frozen backbone and only minimal parameter tuning, CoAR effectively injects the specific subject concept into the AR model and mitigates overfitting through multi-faceted regularization.
Extensive experiments demonstrate the superiority of CoAR in subject fidelity, instruction alignment, and parameter efficiency. Compared to naive DreamBooth adaptation in AR models, CoAR achieves significant performance gains while training approximately fewer than 0.1M parameters. Against existing AR-based customization methods, CoAR achieves competitive or superior results while training with no more than 0.15% of the parameters used by other approaches. For stylization tasks, our training-free method achieves subject identity preservation and high-fidelity stylization across diverse scenes, matching the modern LoRA-based competitor (Shah et al. 2024) that requires additional training. Our main contributions are summarized as follows:
  • We propose a novel framework called CoAR for AR customization. Unlike mainstream LoRA-based approaches, CoAR injects subject concepts into minimal context tokens without modifying pre-trained parameters.
  • We propose a training-free customization strategy that enables arbitrary subject–style combinations during inference. By concatenating subject and style context tokens, our method achieves high-quality stylization while preserving subject identity and contextual consistency without any additional training. Some generated samples are shown in Fig.˜1
  • Experimental results demonstrate that CoAR achieves competitive or superior performance on subject-driven personalized generation, while requiring substantially fewer trainable parameters. Furthermore, it achieves performance comparable to modern diffusion-based approaches in subject-style personalized generation.
Related Works
Autoregressive Image Generation
In recent years, autoregressive (AR) image generation has rapidly evolved from pixel-level sequential modeling to token-level generation and further to multi-modal unified architectures. Early approaches (Van den Oord et al. 2016; Van Den Oord, Kalchbrenner, and Kavukcuoglu 2016) operated directly in the pixel space, which imposed heavy computational demands for generating high-resolution images. The VQ-VAE family (Ramesh et al. 2021; Van Den Oord, Vinyals et al. 2017) subsequently discretizes images into tokens, followed by training a Transformer (Vaswani et al. 2017) model using the Next Token Prediction (NTP) task to generate images. Recent research has explored two major directions: One is Visual autoregressive (VAR) models (Tian et al. 2024; Han et al. 2025). VAR models adopt a next-scale prediction framework that significantly improves image quality. The other is multi-modal unified AR models (Team 2024; Liu et al. 2024a; Xin et al. 2025). Chameleon (Team 2024) employs an early-fusion strategy to mix text and image tokens into a single sequence, enabling unified text-to-image, captioning, and inpainting generation. Lumina-mGPT (Liu et al. 2024a) further incorporates progressive FP-SFT, substantially enhancing high-resolution image synthesis. Lumina-mGPT 2.0 (Xin et al. 2025) approaches state-of-the-art diffusion models within a pure AR framework while maintaining strong multi-task generalization. These unified multi-modal AR models demonstrate strong potential as a foundation for versatile and scalable vision-language generation systems.
Figure 2: Overview of the proposed CoAR framework for subject-driven personalization in multi-modal autoregressive models. (a) Layerwise Multimodal Context Learning, where learnable context tokens are injected into multiple Transformer layers for concept representation. (b) Dual Prior Preservation (DPP) regularizes the customized distribution against the pre-trained model to mitigate overfitting and language drift. (c) Context-Aware Self-Regularization (CASR) initializes and constrains context tokens towards the subject embedding space to enhance fidelity and re-contextualization. (d) Training-free subject–style composition by directly combining subject and style tokens to enable flexible customized generation.
Subject-Driven Image Generation
Subject-driven image generation aims to capture the concept of a specific subject from a limited set of reference images and generalize it to novel contexts. Existing approaches can broadly be categorized into diffusion-based and AR-based methods. Diffusion-based approaches (Ruiz et al. 2023; Chen et al. 2023a; Han et al. 2023; Kumari et al. 2023; Voynov et al. 2023; Gal et al. 2022) typically adapt the distribution of a pre-trained model to that of the target subject via either full fine-tuning or low-rank adaptation (LoRA) (Hu et al. 2022). During inference, a special identifier is employed to sample from the adapted distribution, allowing for customized generation. As the potential of AR models in image generation continues to be exploited, recent studies (Chung et al. 2025; Zheng et al. 2025; Sun et al. 2025; Wu et al. 2025) have begun to investigate subject-driven customization specifically tailored for AR models. For VAR-based models, ARBooth (Chung et al. 2025) introduces selective layer tuning to reduce complexity and employs prior distillation to mitigate the issue of language drift. FreeLoRA (Zheng et al. 2025) proposes a training-free multi-subject customization approach for VAR models. For multimodal AR models, PersonalAR (Sun et al. 2025) validates the combination of DreamBooth (Ruiz et al. 2023) and TI (Gal et al. 2022) within a LoRA framework on Lumina-mGPT. Proxy-Tuning (Wu et al. 2025) further enhances capability in subject-driven image generation by leveraging diffusion models as proxies.
These approaches follow the diffusion-based customization paradigm, adapting the generation backbone to specific subject distributions. A concurrent work, Yo’Chameleon (Nguyen et al. 2025), also introduces context optimization for multimodal generation of the auto-regressive model but suffers from overfitting and language drift, leading to reduced quality and subject fidelity. In contrast, CoAR represents subject or style information using a compact set of learnable context tokens. This eliminates the need to update backbone parameters while preserving the model’s generative capacity. By decoupling customization from backbone adaptation, CoAR offers an effective alternative to prior approaches.
Methodology
In this work, we propose CoAR, a simple yet effective framework for subject-driven customization in multi-modal autoregressive (AR) models. The overall CoAR framework is shown in Fig. 2, consisting of:
  • Layer-wise Multi-modal Contextual Learning: injecting learnable multimodal tokens into multiple Transformer layers.
  • Dual Prior Preservation: regularizing training with the frozen pre-trained model.
  • Context-Aware Self-Regularization: initializing and constraining context tokens with subject embeddings to prevent overfitting.
Preliminary
We first introduce the AR: in multimodal autoregressive (AR) text-to-image generation, both text and images are represented as sequences of discrete tokens drawn from a shared vocabulary $\mathcal{V}$. A text prompt is encoded as:
$\mathbf{y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L}\},\quad\mathbf{y}_{l}\in\mathcal{V},$ (1)
and the target image is represented as
$\mathbf{x}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\},\quad\mathbf{x}_{t}\in\mathcal{V},$ (2)
where $L$ and $T$ denote the lengths of the text and image token sequences, respectively.
AR model generates the image autoregressively by iteratively sampling the next token conditioned on the text and all previously generated tokens:
$\mathbf{x}_{t}\sim p_{\theta}(\mathbf{x}_{t}\mid\mathbf{x}_{<t},\mathbf{y}),\quad t=1,\dots,T,$ (3)
where $\mathbf{x}_{<t}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{t-1}\}$ and $\theta$ are the model parameters.
Training follows the next-token prediction paradigm, minimizing the negative log-likelihood of the ground-truth token sequence conditioned on the text prompt:
$\mathcal{L}_{\text{NTP}}=-\sum_{t=1}^{T}\log p_{\theta}(\mathbf{x}_{t}\mid\mathbf{x}_{<t},\mathbf{y}),$ (4)
where $\mathbf{x}_{t}$ denotes the token at time step $t$, $\mathbf{x}_{<t}$ represents the preceding tokens, and $y$ is the input text prompt. This objective is implemented using the standard cross-entropy loss over the shared vocabulary $\mathcal{V}$.
Layerwise Multimodal Context Learning Strategy
We introduce a Layerwise Multimodal Contextual Learning (LMCL) strategy to capture the target subject concept while keeping the pre-trained model entirely frozen, requiring only a small number of learnable parameters for effective customization.
Specifically, for all images from the same subject, we define a shared set of multimodal learnable tokens as the representation of the subject concept:
$\mathbf{P}=\{\mathbf{p}_{{[v]}},\mathbf{p}_{{I}}\},$ (5)
where $\mathbf{p}_{{[v]}}\in\mathbb{R}^{1\times N\times D}$ denotes the learnable tokens for the text modality with the number of insertion layers $N$ and the token embedding dimension $D$. Similarly, $\mathbf{p}_{{I}}$ represents the learnable tokens for the image modality with the same shape $1\times N\times D$.
For each subject image, we construct a textual template using a shared prompt a photo of [V] [Class], where [V] is a placeholder representing the subject identity and [Class] is the class name. Given the original token sequence at the $i_{th}$ Transformer layer $\mathbf{U}_{i}$ (omitting special start/end tokens for simplicity):
$\mathbf{U}_{i}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}.$ (6)
As shown in Fig.˜2(a), we insert the learnable tokens as:
$\mathbf{U}_{i}^{\prime}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{p}_{{[v]}}^{(i)},\dots,\mathbf{y}_{L},\mathbf{p}_{{I}}^{(i)},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\},$ (7)
where $\mathbf{p}_{{[v]}}^{(i)}$ replaces the placeholder [V] in the text sequence, and $\mathbf{p}_{{I}}^{(i)}$ is inserted before the first image token.
Finally, we optimize the representation of these multimodal tokens using Eq. (4) to progressively inject the concept of the specific subject into them. By learning only $\{\mathbf{p}_{[v]},\mathbf{p}_{I}\}$ while keeping the backbone frozen, LMCL achieves efficient subject-driven customization with minimal parameter overhead.
Dual Prior Preservation
To mitigate language drift and encourage generation diversity, we propose Dual Prior Preservation (DPP) to make the prior preservation introduced in DreamBooth (Ruiz et al. 2023) compatible with training of the multimodal AR model. We first generate a small set (6–8) of class images using the pre-trained model with the prompt “a photo of a [Class]”. These generated images, along with their corresponding text prompts, are tokenized into an input sequence $\mathbf{U}_{{cls}}$. We then perform a dual forward pass on $\mathbf{U}_{{cls}}$:
  • Forwarding the original $\mathbf{U}_{{cls}}$ through the pretrained model to obtain a zero-shot output distribution ${\text{logits}}_{{zs}}$.
  • Inserting layerwise image context tokens into $\mathbf{U}_{{cls}}$ yields $\mathbf{U}_{{cls}}^{\prime}$ and forward through the current model to obtain $\text{logits}_{{prior}}$.
The DPP loss is computed as:
$\mathcal{L}_{{DPP}}$ $=\lambda_{1}\cdot\mathcal{L}_{{NTP_{cls}}}({\text{logits}}_{{prior}},\text{labels}_{cls})$ (8)
$+\lambda_{2}\cdot D_{{KL}}({\rm{logits}}_{{zs}}\parallel{\rm{logits}}_{{prior}}),$
where $\mathcal{L}_{{NTP_{cls}}}$ denotes the Next Token Prediction (NTP) loss on the class images in Eq. (4), $\text{labels}_{cls}$ are the ground-truth visual tokens of the class images, $D_{{KL}}$ is the Kullback–Leibler (KL) divergence measuring the distance between the two output distributions, which is formulated as:
$\begin{aligned} D_{{KL}}({\rm{logits}}_{{zs}}\parallel{\rm{logits}}_{{prior}})&=\\ \sum_{v\in\mathcal{V}}{\rm{logits}}_{{zs}}(v)\log&\frac{{\rm{logits}}_{{zs}}(v)}{{\rm{logits}}_{{prior}}(v)}\end{aligned},$ (9)
and $\lambda_{1},\lambda_{2}$ are coefficients balancing the NTP and KL regularization terms.
The dual-path design extends DreamBooth’s prior preservation from continuous diffusion latents to discrete token modeling, making it compatible with multi-modal AR frameworks.
Context-Aware Self-Regularization
To address overfitting and improve re-contextualization in subject-driven customization, we propose the Context-Aware Self-Regularization (CASR). As shown in Fig.˜2(c), given the reference images of the specific subject, we feed all samples into the frozen pre-trained model and compute the average embedding of all image tokens in each layer $i$, denoted as $\mathbf{E}_{{subject}}^{(i)}$. We initialize the learnable image context token $\mathbf{p}_{I}^{(i)}$ with $\mathbf{E}_{{subject}}^{(i)}$, enabling training to start from a representation closer to the subject’s distribution and accelerating convergence. The CASR loss is defined as:
$\mathcal{L}_{{CASR}}=\frac{1}{N}\sum_{i=1}^{N}\big{\|}\mathbf{p}_{I}^{(i)}-\mathbf{E}_{{subject}}^{(i)}\big{\|}^{2},$ (10)
where $\|\cdot\|^{2}$ denotes the squared $\ell_{2}$-norm measuring the Euclidean distance and $N$ is the total number of insertion layers.
CASR regularizes the context tokens by constraining them towards the subject’s embedding space, effectively preventing overfitting and significantly improving re-contextualization performance.
Overall Objective
The final training objective combines the autoregressive next-token prediction with the proposed regularization terms:
$\mathcal{L}_{obj}=\mathcal{L}_{{NTP}}+\alpha\mathcal{L}_{{DPP}}+\beta\mathcal{L}_{{CASR}},$ (11)
where $\mathcal{L}_{{NTP}}$ denotes the next-token prediction loss on the subject images, $\alpha$ and $\beta$ are two factors that balance the contribution of the DPP and CASR losses, respectively.
Method DINO CLIP-I CLIP-T Trainable Params Training Stage
Real Imges 0.774 0.885 N/A N/A N/A
TI (Ruiz et al. 2023) 0.569 0.780 0.255 N/A
DreamBooth(SD) (Ruiz et al. 2023) 0.668 0.803 0.305 N/A
DreamBooth (Imagen) (Ruiz et al. 2023) 0.696 0.812 0.306 N/A One-Stage
DreamBooth (Lumina-mGPT) (Wu et al. 2025) 0.5088 0.6752 0.2956 N/A
DreamBooth (Lumina-mGPT)+LoRA (Wu et al. 2025) 0.5338 0.6974 0.2652 N/A
PersonalAR (Lumina-mGPT)+LoRA (Sun et al. 2025) 0.657 0.778 0.312 50.4M
PersonalAR (Lumina-mGPT) (Sun et al. 2025) 0.671 0.805 0.302 1610.6M Two-Stage
Proxy-Tuning (Lumina-mGPT/SD3.5)+LoRA (Wu et al. 2025) 0.7523 0.8091 0.3123 142.6M
CoAR (Lumina-mGPT) (Ours) 0.7226 0.8151 0.3184 0.073M One-Stage
Table 1: Performance comparison of our CoAR and the recent competitors on the DreamBench dataset (Ruiz et al. 2023). The best results are marked in bold.
Figure 3: Qualitative comparison of subject-driven personalization on the DreamBench benchmark, selecting PersonalAR and Proxy-Tuning for comparison. The results for both competitors are taken directly from their original publications.
Customize Any Subject in Any Style
Since CoAR leaves the pre-trained AR model parameters untouched, it fully leverages the strong multimodal understanding and generative capacity of autoregressive models, which naturally support zero-shot subject–style composition. This design enables training-free customization of arbitrary subjects in arbitrary styles simply by concatenating their respective context tokens, allowing for flexible, efficient, and scalable rendering without fine-tuning. To further ensure clean and controllable generation, we introduce an Identity Mask during inference, as shown in Fig.˜2(d), which explicitly restricts attention flow between subject and style tokens. This mechanism prevents semantic interference and ensures that each token type contributes distinctly and independently to the final generation.
Experiments
Experimental Details
Figure 4: Re-contextualization with our CoAR on DreamBench (Ruiz et al. 2023) and StyleDrop (Sohn et al. 2023) datasets.
Benchmark. We evaluate subject-driven customization using the DreamBench dataset (Ruiz et al. 2023), which contains 30 subjects (21 objects, 9 animals), each with 3–5 reference images and 25 test prompts. Objects are tested with 20 recontextualization and 5 property modification prompts, while animals use 10 recontextualization, 10 accessorization, and 5 property modification prompts. For stylization, we adopt the StyleDrop dataset (Sohn et al. 2023) and follow ZipLoRA (Shah et al. 2024), training with a single reference image per style. Additionally, we evaluate subject-style hybrid generation by combining DreamBench subjects with StyleDrop styles.
Implementation Details. All experiments are run on a single H800 GPU using Lumina-mGPT-7B FP-SFT (Liu et al. 2024a) as the backbone. For subject-driven personalization, we insert one text and one image token per layer into the first 9 layers. For stylization with a single reference image, the same token configuration is applied to the first 3 layers. The model is trained with a learning rate of 1e-2 and batch size 1, keeping all other Lumina-mGPT hyperparameters unchanged. We set $\alpha$ = 1e-2, $\beta$ = 5e-4, and both $\lambda_{1}$ and $\lambda_{2}$ to 0.5. Each subject and style is trained for 1000 and 600 steps, respectively.
CoAR (Ours) B-LoRA ZipLoRA
Subject-alignment 0.604 0.579 0.655
Style-alignment 0.594 0.505 0.597
Text-alignment 0.308 0.258 0.272
Table 2: Comparison results of our CoAR and two competitors on the style personalization task.
Evaluation Metrics. For subject-driven customization tasks, we follow established protocols and employ CLIP-I (Radford et al. 2021) and DINO (Caron et al. 2021) for subject fidelity evaluation, along with CLIP-T (Radford et al. 2021) for prompt fidelity assessment. Specifically, CLIP-I computes the average cosine similarity between CLIP embeddings of real and generated images. DINO measures the average cosine similarity of ViT-S/16 DINO embeddings between real and generated images. CLIP-T evaluates the cosine similarity between CLIP text embeddings of prompts and CLIP image embeddings of generated outputs. For stylization tasks, we adopt metrics from ZipLoRA (Shah et al. 2024), which use CLIP features for both style and text alignment, and DINO features for subject alignment.
Competitors. For the subject-driven personalization, we compare our approach with several baselines: Textual Inversion (TI) (Gal et al. 2022), DreamBooth (Ruiz et al. 2023), and two DreamBooth-based competitors (Wu et al. 2025), which share the same autoregressive backbone as our method. We also include two recent approaches specifically designed for multimodal autoregressive models (Sun et al. 2025; Wu et al. 2025). For the style personalization, we compare with recent diffusion-based methods: ZipLoRA (Shah et al. 2024) and B-LoRA (Frenkel et al. 2024).
Main Results
Quantitative Results of Subject-Driven Personalization.
Table 1 shows the comparison results of our CoAR and the competitors on the DreamBench dataset. From the results, we have the following observations. First, CoAR achieves superior performance across multiple dimensions. In the CLIP-I metric, CoAR achieves a leading score of 0.8151, surpassing the second-best DreamBooth (Imagen) by 0.4%, demonstrating its exceptional image fidelity. For CLIP-T, CoAR sets a new benchmark at 0.3184, outperforming the second-best method, Proxy-Tuning by 0.29%, highlighting its superior cross-modal alignment capability. While CoAR ranks second on the DINO metric, it is substantially more efficient and user-friendly than the top-performing method, Proxy-Tuning, which requires training an additional diffusion model and generating hundreds to a thousand images per subject for data augmentation.
Quantitative Results of Style Personalization.
Table 2 presents a comparison between our CoAR model and two baselines: B-LoRA (Frenkel et al. 2024) and ZipLoRA (SDXL) (Shah et al. 2024). The results demonstrate that CoAR outperforms B-LoRA across all three evaluation metrics, achieving the highest text-alignment score among all competitors. While CoAR trails ZipLoRA slightly in subject alignment (0.604 vs. 0.655), it achieves comparable style preservation (0.594 vs. 0.597), superior text alignment (0.308 vs. 0.272), and does so without requiring any additional training.
Qualitative Results of Subject-Driven Personalization.
Fig.˜3 illustrates some generated samples of our CoAR compared with two most related competitors, i.e., PersonalAR (Sun et al. 2025) and Proxy-Tuning (Wu et al. 2025). From the results, we observe that images generated by CoAR exhibit superior prompt fidelity. For instance, given the instruction “wearing a black top hat and a monocle", neither PersonAR nor Proxy-Tuning successfully generates the “monocle", whereas our CoAR does. Moreover, our generated images better adhere to physical plausibility. In the second row of PersonAR, the generated vase and backpack are both floating, which is unrealistic, while such artifacts are absent in our results. More qualitative results are provided in the supplementary materials.
Figure 5: Qualitative comparison of our CoAR and LoRA-based method for style personalization tasks.
Qualitative Results of Style Personalization.
As shown in Fig.˜5, CoAR demonstrates superior performance compared to B-LoRA. While ZipLoRA preserves more subject details, such as color, CoAR exhibits stronger adherence to instructions when generating in novel scenes (e.g., “in the jungle” or “floating in water”). Additional qualitative results on style customization are presented in Fig.˜4, showing that CoAR effectively preserves the given style when generalizing the reference subject to novel scenes.
Ablation Studies
Impacts of Different Losses. To assess the contribution of each loss component in our framework, we perform an ablation study by incrementally applying the LMCL, DPP, and CASR losses and evaluating performance on DINO, CLIP-I, and CLIP-T metrics, as shown in Tab. 3. Only applying LMCL yields baseline scores of 0.6610 (DINO), 0.7647 (CLIP-I), and 0.3096 (CLIP-T). Introducing DPP leads to consistent improvements across all metrics, increasing DINO to 0.6902 and CLIP-I to 0.7919. Although CLIP-T slightly decreases to 0.2968, the overall alignment benefits are clear. When CASR is further added, we observe the best overall performance: 0.7226 (DINO), 0.8151 (CLIP-I), and 0.3184 (CLIP-T), confirming that each loss component contributes positively to the alignment quality and subject fidelity.
LMCL DPP CASR DINO CLIP-I CLIP-T
0.6610 0.7647 0.3096
0.6902 0.7919 0.2968
0.7226 0.8151 0.3184
Table 3: Impacts of different losses on the subject-driven personalization task.
Impacts of Layer Numbers. To evaluate the impacts of depth of multimodal context token insertion, we insert tokens at layers 1, 3, 9, and 24. As illustrated in Fig.˜6. CLIP-T remains relatively stable across depths, while CLIP-I and DINO follow a rise-then-fall trend. Notably, inserting tokens at deeper layers (e.g., layer 24) leads to a decline in subject fidelity. We attribute this to overfitting caused by limited reference data, which becomes more pronounced as token influence shifts to higher-level semantic layers.
More experimental results are provided in the appendix.
Figure 6: Impacts of insertion depth of multimodal context tokens on performance.
Conclusion
In this paper, we introduced CoAR, a novel and efficient framework for subject-driven and style-aware customization in multimodal autoregressive (AR) models. Unlike conventional fine-tuning or LoRA-based approaches, CoAR injects subject concepts directly into the pretrained distribution through a small number of learnable context tokens, avoiding any modification to pretrained parameters. Extensive experiments across personalization and stylization tasks demonstrate that CoAR achieves superior subject fidelity, instruction alignment, and style preservation while training fewer than 0.1M parameters, significantly more parameter-efficient than existing approaches. Moreover, our training-free strategy matches or surpasses modern LoRA-based diffusion methods, highlighting the potential of multimodal AR models for scalable and flexible personalization.
Appendix
Appendix A Training and Evaluation Details
Training Details. The training details of CoAR on different datasets are shown in Table˜4. For the style personalization task, only the provided single reference image is utilized. Consequently, the hyperparameters for Dual Prior Preservation (DPP) are not applied.
Evaluation Details. For the subject-style generation task, we selected 8 subjects from DreamBench (Ruiz et al. 2023), including 4 animals and 4 objects, as well as 6 styles from StyleDrop (Sohn et al. 2023). These yielded a total of 48 unique subject-style combinations, on which we evaluated our CoAR, ZipLoRA (Shah et al. 2024), and B-LoRA (Frenkel et al. 2024). For each combination, we assigned 10 distinct recontextualization prompts for both objects and living subjects, as illustrated in Table˜5.
DreamBench StyleDrop
lr 1e-2 1e-2
batch size 1 1
resolution 768 768
$\lambda_{1}$ 0.5 N/A
$\lambda_{2}$ 0.5 N/A
$\alpha$ 1e-2 N/A
$\beta$ 5e-4 5e-4
Training Steps 1000 600
Context token layers 9 3
Table 4: Training settings of CoAR on DreamBench(Ruiz et al. 2023) and StyleDrop (Sohn et al. 2023) datasets.
subject context
objects in the snow
on the beach
with a city in the background
on top of a dirt road
on top of green grass
in the jungle
on the mountain
floating in water
on a picnic table
on top of a wooden floor
living subjects wearing a hat
with a crown
riding a bicycle
sleeping
in a boat
in the jungle
on the mountain
floating in water
on a picnic table
playing with a ball
Table 5: Prompt templates categorized by subject type.
Appendix B Details of Parameter Calculation
For PersonalAR (Sun et al. 2025), we directly reference the parameter count reported in its original paper. As for Proxy-Tuning (Wu et al. 2025), since the paper does not provide explicit parameter statistics and the code is not publicly available, we provide a coarse estimate based on the hyperparameter configurations described in the paper.
Proxy-Tuning is a two-stage approach where a diffusion model is first trained, followed by an autoregressive (AR) model that is trained on data generated by the diffusion model. Both models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al. 2022). In our estimation, the diffusion model is based on Stable Diffusion 3.5 (SD3.5 Large) (Esser et al. 2024), while the AR model is Lumina-mGPT 7B (Liu et al. 2024a). We estimate the total number of trainable LoRA parameters based on the typical LoRA application settings.
LoRA Parameter Calculation
For a weight matrix $W\in\mathbb{R}^{d\times d}$, LoRA introduces two trainable matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times d}$, where $r$ is the LoRA rank. The number of trainable parameters for each projection is:
$\text{LoRA params}=2dr$
We assume LoRA is applied to all attention projections: $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$.
Diffusion Model (SD3.5)
SD3.5 is based on the DiT (Peebles and Xie 2023) architecture, with 24 Transformer (Vaswani et al. 2017) blocks. Each block includes one self-attention and one cross-attention module. Assuming a hidden size $d=3072$, the LoRA (rank=64) parameters per attention module are:
$4\times 2\times d\times r$ $=8dr$ (12)
[…]
There are a total of $24\times 2=48$ attention modules, leading to:
[…]
AR Model (Lumina-mGPT 7B)
For the AR model, we assume a hidden size $d=4096$ and 32 Transformer layers. Applying LoRA (rank = 64) to the four attention projections in each layer results in:
$4\times 2\times d\times r$ $=$ (14)
$=8dr$
[…]
leading to:
[…]
Total Parameter Count
Summing both components, the total number of trainable LoRA parameters under Proxy-Tuning is approximately:
[…]
This estimate assumes LoRA is applied to all attention projections and excludes feed-forward layers. Including FFNs would increase the total parameter count further, but standard LoRA configurations typically restrict adaptation to attention modules only.
Appendix C Additional Experiments
Ablation
As shown in Fig.˜7, we first analyze the results using only Layer-wise Multi-modal Contextual Learning (LMCL), and then with the addition of the Dual Prior Preservation (DPP) loss. With the DPP constraint, the model better preserves the subject’s identity. However, when recontextualizing the subject in the scene “on a cobblestone street”, it still struggles to accurately render the background. In contrast, incorporating Context-Aware Self-Regularization (CASR) enables the model to successfully generalize to the new scene.
Fig.˜8 provides a visual illustration of the effect of our proposed CASR loss. When the loss weight is too small, the model tends to overfit: although subject fidelity remains high, generalization to novel contexts is limited. Conversely, when the weight is too large, the context tokens become under-optimized, leading to a significant degradation in subject fidelity.
Figure 7: Visualization of the effects of different losses on the generated samples.
Figure 8: Visualization of CASR.
Figure 9: Qualitative results of subject-driven customization with CoAR.
Figure 10: Qualitative results of subject-driven customization with CoAR.
Additional Qualitative Results
Figs.˜9, 10, 12, 11 and 13 presents additional qualitative results of subject-driven personalization and style personalization.
Figs.˜9 and 10 demonstrate that given only a few reference images, CoAR successfully performs subject-driven generation across various tasks, including recontextualization (e.g., novel backgrounds), property modification (e.g., color and shape), and accessorization (e.g., adding glasses or outfits). The results demonstrate strong subject fidelity, adaptability to novel contexts, and fine-grained controllability.
Fig.˜12 presents examples of image generation conditioned on different artistic styles. Each row represents a different subject, while each column corresponds to a specific style. The generated images highlight the model’s ability to retain subject identity while adapting visual attributes such as texture, color palette, and rendering technique to match the target style.
Fig.˜13 illustrates the effects of diverse contextual prompts on image generation for various subjects across different visual styles. Each row corresponds to a distinct subject, while each column varies the context.
Figure 11: CoAR demonstrates effective style personalization conditioned on only a single reference image.
Figure 12: Style-conditioned image generation across subjects without contextual prompts.
Figure 13: Impact of contextual prompts on image generation across subjects and styles.

### v2

@@
-CoAR: Concept Injection into Autoregressive Models for Personalized Text-to-Image Generation
+DCoAR: Deep Concept Injection into Unified Autoregressive Models for Personalized Text-to-Image Generation
 Abstract
-The unified autoregressive (AR) model excels at multimodal understanding and generation, but its potential for customized image generation remains underexplored.
-Existing customized generation methods rely on full fine-tuning or adapters, making them costly and prone to overfitting or catastrophic forgetting.
-In this paper, we propose CoAR, a novel framework for injecting subject concepts into the unified AR models while keeping all pre-trained parameters completely frozen.
-CoAR learns effective, specific subject representations with only a minimal number of parameters using a Layerwise Multimodal Context Learning strategy.
-To address overfitting and language drift, we further introduce regularization that preserves the pre-trained distribution and anchors context tokens to improve subject fidelity and re-contextualization.
-Additionally, CoAR supports training-free subject customization in a user-provided style.
-Experiments demonstrate that CoAR achieves superior performance on both subject-driven personalization and style personalization, while delivering significant gains in computational and memory efficiency.
-Notably, CoAR tunes less than 0.05% of the parameters while achieving competitive performance compared to recent Proxy-Tuning.
-Code: https://github.com/KZF-kzf/CoAR.
-Figure 1: Visualization of samples generated by our proposed CoAR, showcasing subject personalization (left), where a few reference images are provided, and style personalization (right), where the subject is rendered in user-specified artistic styles without additional training.
-Introduction
-Subject-driven image generation (Chen et al.
-2023b; Ruiz et al.
-2023; Li, Li, and Hoi 2023; Shin et al.
-2025) aims to generate images of a specific subject or concept in new scenes, given only a few reference images.
-This capability has found applications in virtual try-on (Ye et al.
-2025; Zhang et al.
-2024), personalized content creation (Liu et al.
-2024b), and advertising design (Chen et al.
-2023b).
-Most of the existing approaches, such as DreamBooth (Ruiz et al.
-2023) and its extensions (Chen et al.
-2023a; Han et al.
-2023; Kumari et al.
-2023), customize generation by adapting the distribution of pre-trained diffusion models to match the subject distribution through full fine-tuning or parameter-efficient techniques like LoRA (Hu et al.
-2022).
-These methods typically associate the subject with a unique identifier [V], enabling personalized generation via prompt injection during inference.
-However, such techniques are predominantly designed for diffusion models (Esser et al.
-2024; Rombach et al.
-2022; Podell et al.
-2023) and are not directly transferable to multimodal autoregressive (AR) models.
-Applying these methods naively to AR models poses two challenges.
-  • Risk of overfitting.
-AR models are pretrained on large-scale multimodal corpora and expected to support both understanding and generation.
-In contrast, the target subject typically has only a few reference images (3–5).
-Such scarce data makes it extremely difficult to adapt the pre-trained models to the subject distribution, resulting in a risk of overfitting and catastrophic forgetting of the pre-trained knowledge.
-  • Inefficiency. Extensively adjusting the pretrained model parameters to learn a single concept with the limited number of images per subject is intuitively inefficient.
-We argue that instead of fine-tuning the pretrained model on limited data to shift its distribution toward a target subject, it is more effective to inject subject concepts directly into the model’s existing distribution, thereby fully leveraging the generative capabilities of large models.
-To this end, we introduce Concept Injection for AR Models (CoAR), an extremely efficient framework for customized generation in multimodal AR models.
-CoAR adopts a Layer-wise Multimodal Context Learning (LMCL) strategy, encoding subject-specific information through a small number of learnable prompts without modifying any pretrained parameters.
-To guide learning while preserving the model’s original capabilities, we propose a Dual Prior Preservation (DPP) loss that constrains training.
-Additionally, we design a Context-Aware Self-Regularization (CASR) loss to improve re-contextualization stability and output quality.
-Built upon this foundation, CoAR also supports training-free customization, enabling arbitrary subject–style combinations during inference while maintaining strong re-contextualization performance.
-With a frozen backbone and only minimal parameter tuning, CoAR effectively injects the specific subject concept into the AR model and mitigates overfitting through multi-faceted regularization.
-Extensive experiments demonstrate the superiority of CoAR in subject fidelity, instruction alignment, and parameter efficiency.
-Compared to naive DreamBooth adaptation in AR models, CoAR achieves significant performance gains while training approximately fewer than 0.1M parameters.
-Against existing AR-based customization methods, CoAR achieves competitive or superior results while training with no more than 0.15% of the parameters used by other approaches.
-For stylization tasks, our training-free method achieves subject identity preservation and high-fidelity stylization across diverse scenes, matching the modern LoRA-based competitor (Shah et al.
-2024) that requires additional training.
+The unified autoregressive (AR) model excels at multimodal understanding and generation.
+However, its full potential in the domain of customized image generation has yet to be fully realized.
+Existing customization approaches for unified AR models face a fundamental dilemma: adaptation-based methods suffer from overfitting and scalability bottlenecks, while concept-injection paradigms are constrained by a shallow injection strategy that leads to poor visual fidelity and impaired re-contextualization.
+To address this, we propose DCoAR, a novel deep concept injection framework that maintains a completely frozen pre-trained model.
+DCoAR deeply integrates new concepts through a Layer-wise Multimodal Context Learning (LMCL) strategy, which is stabilized by a multi-faceted regularization scheme: a Dual Prior Preservation (DPP) loss to mitigate semantic drift and a Context-Aware Self-Regularization (CASR) loss to enhance re-contextualization.
+The framework also enables training-free subject customization in user-provided styles.
+Experiments demonstrate that DCoAR significantly outperforms previous injection-based methods and achieves performance competitive with adaptation-based approaches while requiring substantially fewer trainable parameters.
+Figure 1: Visualization of samples generated by our proposed DCoAR, showcasing subject personalization (left), where a few reference images are provided, and style personalization (right), where the subject is rendered in user-specified artistic styles without additional training.
+1 Introduction
+Unified multimodal autoregressive (AR) models [team2024chameleon, lumina_mgpt, mmr1, stylear, editar, acdit], have established a powerful paradigm for processing interleaved text and image data, demonstrating remarkable success across a range of generative and understanding tasks.
+Despite their success, the application of these models to personalized image generation has received comparatively less attention, leaving their full potential in this specific task open for exploration.
+Personalized image generation[chen2023subject, ruiz2023dreambooth, li2023blip, shin2025large, liu2024llm4genleveragingsemanticrepresentation, mosaicmultisubjectpersonalizedgeneration, TFcustom, she2025customvideox3dreferenceattention] aims to generate images of a specific subject or concept in new scenes, given only a few reference images.
+This capability has found applications in virtual try-on [ye2025slot, zhang2024two], personalized content creation [liu2024make], and advertising design [chen2023subject].
+Personalization for unified AR models remains fundamentally challenging, and existing approaches fall short in ways that are increasingly incompatible with real-world demands.
+On the one hand, fine-tuning–based methods, inspired by DreamBooth [ruiz2023dreambooth, chen2023disenbooth, han2023svdiff, customdiff, OmniControl, EasyControl] and implemented through PEFT techniques such as LoRA [Prefixtuning, coop, hu2022lora, personalAR, wu2025proxytuning], attempt to rewrite model weights for each new concept.
+Yet AR models are far more brittle than diffusion models: adapting millions of parameters using only 3–5 images almost inevitably causes overfitting, catastrophic forgetting, and distortion of pre-trained priors.
+Worse still, maintaining a separate set of weights for every concept leads to unbounded storage costs, making this strategy fundamentally non-scalable for consumer-facing or large-scale personalization systems.
+On the other hand, token-injection approaches [UniCTokens, yoChameleon] promise scalability by keeping the backbone frozen, but they suffer from a deeper structural flaw.
+Injecting concept information only at the input layer forces the signal to pass through dozens of transformer blocks without reinforcement.
+This shallow injection bottleneck results in severe semantic attenuation, where fine-grained identity cues vanish as depth increases, and the model fails to bind the concept to complex prompts.
+Consequently, these methods struggle with visual fidelity, contextual adaptability, and semantic consistency, falling short of fine-tuning counterparts.
+To overcome the constraints of shallow concept injection, we introduce Deep Concept Injection for unified AR models (DCoAR), an efficient and extensible framework for customized image generation.
+The core of our DCoAR framework is the Layer-wise Multimodal Context Learning (LMCL) strategy that encodes subject-specific information into a compact set of learnable tokens and injects them across multiple transformer layers, enabling deeper semantic alignment and high-fidelity feature propagation without modifying the frozen backbone.
+To ensure the robustness and quality of this deep injection process, DCoAR incorporates a multi-faceted regularization scheme consisting of two components.
+First, a Dual Prior Preservation (DPP) loss constrains the model against semantic drift via constraining the concept-conditioned features to stay close to the model’s native feature distribution while penalizing deviations from its latent feature gradients, ensuring that personalization does not compromise the inherent generative priors of the base AR model.
+Second, a Context-Aware Self-Regularization (CASR) enhances re-contextualization by preventing the concept tokens from over-specializing to the limited training prompts.
+CASR enforces cross-context consistency by aligning the model’s internal representations under both original and perturbed contexts, encouraging the tokens to stay flexible rather than memorizing appearance details.
+It further penalizes representation collapse across layers, ensuring that the injected concept signals remain discriminative and generalizable during deep propagation.
 Our main contributions are summarized as follows:
-  • We propose a novel framework called CoAR for AR customization.
-Unlike mainstream LoRA-based approaches, CoAR injects subject concepts into minimal context tokens without modifying pre-trained parameters.
-  • We propose a training-free customization strategy that enables arbitrary subject–style combinations during inference.
-By concatenating subject and style context tokens, our method achieves high-quality stylization while preserving subject identity and contextual consistency without any additional training.
-Some generated samples are shown in Fig.˜1
-  • Experimental results demonstrate that CoAR achieves competitive or superior performance on subject-driven personalized generation, while requiring substantially fewer trainable parameters.
-Furthermore, it achieves performance comparable to modern diffusion-based approaches in subject-style personalized generation.
-Related Works
-Autoregressive Image Generation
+  • We propose a deep concept injection framework for unified AR models.
+Unlike prior methods constrained by shallow, early-layer conditioning, DCoAR introduces a Layer-wise Multimodal Context Learning (LMCL) strategy that injects concept signals throughout the transformer stack, enabling persistent feature reinforcement and richer cross-modal interaction.
+This design achieves robust concept propagation and high-fidelity personalization without modifying a single parameter of the frozen backbone.
+  • We introduce a training-free subject–style composition strategy.
+By directly concatenating the learned context tokens for a subject and a style, DCoAR enables arbitrary, plug-and-play combinations at inference time, achieving high-quality, context-aware stylization without any additional optimization or fine-tuning.
+Some generated samples are shown in Fig.˜1.
+  • Experimental results show that DCoAR sets a new state of the art for unified AR model personalization.
+Furthermore, DCoAR delivers strong gains in subject fidelity, instruction alignment, and parameter efficiency, requiring fewer than 0.1M trainable parameters.
+For stylization, our training-free pipeline produces high-fidelity results that rival modern LoRA-based methods without any additional training.
+2 Related Works
+2.1 Autoregressive Image Generation
 In recent years, autoregressive (AR) image generation has rapidly evolved from pixel-level sequential modeling to token-level generation and further to multi-modal unified architectures.
-Early approaches (Van den Oord et al.
-2016; Van Den Oord, Kalchbrenner, and Kavukcuoglu 2016) operated directly in the pixel space, which imposed heavy computational demands for generating high-resolution images.
-The VQ-VAE family (Ramesh et al.
-2021; Van Den Oord, Vinyals et al.
-2017) subsequently discretizes images into tokens, followed by training a Transformer (Vaswani et al.
-2017) model using the Next Token Prediction (NTP) task to generate images.
-Recent research has explored two major directions: One is Visual autoregressive (VAR) models (Tian et al.
-2024; Han et al.
-2025).
+Early approaches [pixelcnn, pixelrnn] operated directly in the pixel space, which imposed heavy computational demands for generating high-resolution images.
+The VQ-VAE family [dalle, vqvae] subsequently discretizes images into tokens, followed by training a Transformer [transformer] model using the Next Token Prediction (NTP) task to generate images.
+Recent research has explored two major directions: One is Visual autoregressive (VAR) models [var, infinity].
 VAR models adopt a next-scale prediction framework that significantly improves image quality.
-The other is multi-modal unified AR models (Team 2024; Liu et al.
-2024a; Xin et al.
-2025).
-Chameleon (Team 2024) employs an early-fusion strategy to mix text and image tokens into a single sequence, enabling unified text-to-image, captioning, and inpainting generation.
-Lumina-mGPT (Liu et al.
-2024a) further incorporates progressive FP-SFT, substantially enhancing high-resolution image synthesis.
-Lumina-mGPT 2.0 (Xin et al.
-2025) approaches state-of-the-art diffusion models within a pure AR framework while maintaining strong multi-task generalization.
+The other is multi-modal unified AR models [team2024chameleon, lumina_mgpt, lumina2].
+Chameleon [team2024chameleon] employs an early-fusion strategy to mix text and image tokens into a single sequence, enabling unified text-to-image, captioning, and inpainting generation.
+Lumina‑mGPT [lumina_mgpt] further incorporates progressive FP-SFT, substantially enhancing high-resolution image synthesis.
+Lumina‑mGPT 2.0 [lumina2] approaches state-of-the-art diffusion models within a pure AR framework while maintaining strong multi-task generalization.
 These unified multi-modal AR models demonstrate strong potential as a foundation for versatile and scalable vision-language generation systems.
-Figure 2: Overview of the proposed CoAR framework for subject-driven personalization in multi-modal autoregressive models.
+Figure 2: Overview of the proposed DCoAR framework for subject-driven personalization in multi-modal autoregressive models.
 (a) Layerwise Multimodal Context Learning, where learnable context tokens are injected into multiple Transformer layers for concept representation.
@@
 (d) Training-free subject–style composition by directly combining subject and style tokens to enable flexible customized generation.
-Subject-Driven Image Generation
-Subject-driven image generation aims to capture the concept of a specific subject from a limited set of reference images and generalize it to novel contexts.
-Existing approaches can broadly be categorized into diffusion-based and AR-based methods.
-Diffusion-based approaches (Ruiz et al.
-2023; Chen et al.
-2023a; Han et al.
-2023; Kumari et al.
-2023; Voynov et al.
-2023; Gal et al.
-2022) typically adapt the distribution of a pre-trained model to that of the target subject via either full fine-tuning or low-rank adaptation (LoRA) (Hu et al.
-2022).
-During inference, a special identifier is employed to sample from the adapted distribution, allowing for customized generation.
-As the potential of AR models in image generation continues to be exploited, recent studies (Chung et al.
-2025; Zheng et al.
-2025; Sun et al.
-2025; Wu et al.
-2025) have begun to investigate subject-driven customization specifically tailored for AR models.
-For VAR-based models, ARBooth (Chung et al.
-2025) introduces selective layer tuning to reduce complexity and employs prior distillation to mitigate the issue of language drift.
-FreeLoRA (Zheng et al.
-2025) proposes a training-free multi-subject customization approach for VAR models.
-For multimodal AR models, PersonalAR (Sun et al.
-2025) validates the combination of DreamBooth (Ruiz et al.
-2023) and TI (Gal et al.
-2022) within a LoRA framework on Lumina-mGPT.
-Proxy-Tuning (Wu et al.
-2025) further enhances capability in subject-driven image generation by leveraging diffusion models as proxies.
-These approaches follow the diffusion-based customization paradigm, adapting the generation backbone to specific subject distributions.
-A concurrent work, Yo’Chameleon (Nguyen et al.
-2025), also introduces context optimization for multimodal generation of the auto-regressive model but suffers from overfitting and language drift, leading to reduced quality and subject fidelity.
-In contrast, CoAR represents subject or style information using a compact set of learnable context tokens.
-This eliminates the need to update backbone parameters while preserving the model’s generative capacity.
-By decoupling customization from backbone adaptation, CoAR offers an effective alternative to prior approaches.
-Methodology
-In this work, we propose CoAR, a simple yet effective framework for subject-driven customization in multi-modal autoregressive (AR) models.
-The overall CoAR framework is shown in Fig.
-2, consisting of:
+2.2 Personalized Image Generation
+Personalized image generation aims to capture a specific concept from a few reference images and render it in novel contexts.
+While this task has been extensively studied in diffusion models [ruiz2023dreambooth, chen2023disenbooth, han2023svdiff, customdiff, pplus_TI, textualinversion], its application to unified autoregressive (AR) models presents unique challenges and has led to the emergence of two primary paradigms.
+The first is the adaptation-based paradigm, which employs parameter-efficient fine-tuning (PEFT) like LoRA [hu2022lora] to adapt a model’s weights for a new concept [personalAR, wu2025proxytuning, ARBooth, freelora].
+While capable of high fidelity, this approach is hampered by overfitting and scalability bottlenecks for large-scale deployment.
+The second is the concept-injection paradigm, which inserts concepts into a frozen model via learnable tokens to ensure scalability.
+Pioneering works in this area, such as Yo’Chameleon [yoChameleon] and UniCTokens [UniCTokens], have explored different token designs.
+However, these methods are fundamentally constrained by a shallow injection strategy, where confining the concept to the input layer limits the propagation of visual details, resulting in a fidelity gap compared to adaptation-based approaches.
+In contrast, DCoAR advances this paradigm by introducing a deep injection strategy.
+Through Layer-wise Multimodal Context Learning (LMCL), concept tokens are injected across multiple transformer layers, ensuring robust feature propagation.
+This approach effectively bridges the fidelity gap with adaptation-based methods while preserving the scalability inherent to the concept-injection paradigm.
+3 Method
+In this work, we propose DCoAR, a simple yet effective framework for subject-driven customization in unified autoregressive (AR) models. The overall DCoAR framework is shown in Fig. 2, consisting of:
   • Layer-wise Multi-modal Contextual Learning: injecting learnable multimodal tokens into multiple Transformer layers.
@@
   • Context-Aware Self-Regularization: initializing and constraining context tokens with subject embeddings to prevent overfitting.
-Preliminary
-We first introduce the AR: in multimodal autoregressive (AR) text-to-image generation, both text and images are represented as sequences of discrete tokens drawn from a shared vocabulary $\mathcal{V}$.
-A text prompt is encoded as:
+3.1 Preliminary
+In unified autoregressive (AR) text-to-image generation, both text and images are represented as sequences of discrete tokens drawn from a shared vocabulary $\mathcal{V}$. A text prompt is encoded as:
 $\mathbf{y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L}\},\quad\mathbf{y}_{l}\in\mathcal{V},$ (1)
@@
 This objective is implemented using the standard cross-entropy loss over the shared vocabulary $\mathcal{V}$.
-Layerwise Multimodal Context Learning Strategy
+3.2 Layerwise Multimodal Context Learning Strategy
 We introduce a Layerwise Multimodal Contextual Learning (LMCL) strategy to capture the target subject concept while keeping the pre-trained model entirely frozen, requiring only a small number of learnable parameters for effective customization.
@@
 As shown in Fig.˜2(a), we insert the learnable tokens as:
-$\mathbf{U}_{i}^{\prime}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{p}_{{[v]}}^{(i)},\dots,\mathbf{y}_{L},\mathbf{p}_{{I}}^{(i)},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\},$ (7)
+$\mathbf{U}_{i}^{\prime}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{p}_{[v]}^{(i)},\dots,\mathbf{y}_{L},\mathbf{p}_{I}^{(i)},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\},$
 where $\mathbf{p}_{{[v]}}^{(i)}$ replaces the placeholder [V] in the text sequence, and $\mathbf{p}_{{I}}^{(i)}$ is inserted before the first image token.
@@
 By learning only $\{\mathbf{p}_{[v]},\mathbf{p}_{I}\}$ while keeping the backbone frozen, LMCL achieves efficient subject-driven customization with minimal parameter overhead.
-Dual Prior Preservation
-To mitigate language drift and encourage generation diversity, we propose Dual Prior Preservation (DPP) to make the prior preservation introduced in DreamBooth (Ruiz et al.
-2023) compatible with training of the multimodal AR model.
+3.3 Dual Prior Preservation
+To mitigate language drift and encourage generation diversity, we propose Dual Prior Preservation (DPP) to make the prior preservation introduced in DreamBooth [ruiz2023dreambooth] compatible with training of the multimodal AR model.
 We first generate a small set (6–8) of class images using the pre-trained model with the prompt “a photo of a [Class]”.
@@
 The dual-path design extends DreamBooth’s prior preservation from continuous diffusion latents to discrete token modeling, making it compatible with multi-modal AR frameworks.
-Context-Aware Self-Regularization
+3.4 Context-Aware Self-Regularization
 To address overfitting and improve re-contextualization in subject-driven customization, we propose the Context-Aware Self-Regularization (CASR).
@@
 CASR regularizes the context tokens by constraining them towards the subject’s embedding space, effectively preventing overfitting and significantly improving re-contextualization performance.
-Overall Objective
+3.5 Overall Objective
 The final training objective combines the autoregressive next-token prediction with the proposed regularization terms:
@@
 where $\mathcal{L}_{{NTP}}$ denotes the next-token prediction loss on the subject images, $\alpha$ and $\beta$ are two factors that balance the contribution of the DPP and CASR losses, respectively.
-Method DINO CLIP-I CLIP-T Trainable Params Training Stage
-Real Imges 0.774 0.885 N/A N/A N/A
-TI (Ruiz et al. 2023) 0.569 0.780 0.255 N/A
-DreamBooth(SD) (Ruiz et al. 2023) 0.668 0.803 0.305 N/A
-DreamBooth (Imagen) (Ruiz et al. 2023) 0.696 0.812 0.306 N/A One-Stage
-DreamBooth (Lumina-mGPT) (Wu et al. 2025) 0.5088 0.6752 0.2956 N/A
-DreamBooth (Lumina-mGPT)+LoRA (Wu et al. 2025) 0.5338 0.6974 0.2652 N/A
-PersonalAR (Lumina-mGPT)+LoRA (Sun et al. 2025) 0.657 0.778 0.312 50.4M
-PersonalAR (Lumina-mGPT) (Sun et al. 2025) 0.671 0.805 0.302 1610.6M Two-Stage
-Proxy-Tuning (Lumina-mGPT/SD3.5)+LoRA (Wu et al. 2025) 0.7523 0.8091 0.3123 142.6M
-CoAR (Lumina-mGPT) (Ours) 0.7226 0.8151 0.3184 0.073M One-Stage
-Table 1: Performance comparison of our CoAR and the recent competitors on the DreamBench dataset (Ruiz et al. 2023). The best results are marked in bold.
-Figure 3: Qualitative comparison of subject-driven personalization on the DreamBench benchmark, selecting PersonalAR and Proxy-Tuning for comparison.
-The results for both competitors are taken directly from their original publications.
-Customize Any Subject in Any Style
-Since CoAR leaves the pre-trained AR model parameters untouched, it fully leverages the strong multimodal understanding and generative capacity of autoregressive models, which naturally support zero-shot subject–style composition.
+Figure 3: Qualitative comparison of subject-driven personalization on the DreamBench benchmark.
+As concept-injection methods generally exhibit lower visual fidelity, we focus the qualitative comparison on the more competitive adaptation-based methods.
+Additional comparisons are available in the supplementary material.
+Method Tyeps Method DINO CLIP-I CLIP-T Trainable Params
+Real Imges 0.774 0.885 N/A N/A
+Diffusion Models TI[textualinversion] 0.569 0.780 0.255 N/A
+DreamBooth(SD)[ruiz2023dreambooth] 0.668 0.803 0.305 N/A
+DreamBooth (Imagen)[ruiz2023dreambooth] 0.696 0.812 0.306 N/A
+DreamBooth (Flux)[wu2025proxytuning] 0.732 0.802 0.314 N/A
+Unified AR Models DreamBooth(Lumina-mGPT)[wu2025proxytuning] 0.509 0.675 0.295 N/A
+DreamBooth(Lumina-mGPT)[wu2025proxytuning] +LoRA 0.534 0.697 0.265 N/A
+Yo’Chameleon[yoChameleon] 0.542 0.795 0.225 0.13M
+UniCTokens[UniCTokens] 0.599 0.782 0.304 0.13M
+PersonalAR[personalAR] 0.671 0.805 0.302 1610.6M
+PersonalAR[personalAR] +LoRA 0.657 0.778 0.312 50.4M
+Proxy-Tuning[wu2025proxytuning] 0.752 0.809 0.312 142.6M
+DCoAR(Ours) 0.723 0.815 0.318 0.073M
+Table 1: Performance comparison with state-of-the-art methods on the DreamBench dataset.
+The results for competing methods are directly cited from previous publications and details on parameter count calculation can be found in the supplementary material.
+3.6 Customize Any Subject in Any Style
+Since DCoAR leaves the pre-trained AR model parameters untouched, it fully leverages the strong multimodal understanding and generative capacity of autoregressive models, which naturally support zero-shot subject–style composition.
 This design enables training-free customization of arbitrary subjects in arbitrary styles simply by concatenating their respective context tokens, allowing for flexible, efficient, and scalable rendering without fine-tuning.
@@
 This mechanism prevents semantic interference and ensures that each token type contributes distinctly and independently to the final generation.
-Experiments
-Experimental Details
-Figure 4: Re-contextualization with our CoAR on DreamBench (Ruiz et al. 2023) and StyleDrop (Sohn et al. 2023) datasets.
+4 Experiments
+4.1 Experimental Details
+Figure 4: Training-free subject-style generation with our DCoAR on DreamBench [ruiz2023dreambooth] and StyleDrop [sohn2023styledrop] datasets.
+Figure 5: Re-contextualization with our DCoAR on DreamBench [ruiz2023dreambooth] and StyleDrop [sohn2023styledrop] datasets.
 Benchmark.
-We evaluate subject-driven customization using the DreamBench dataset (Ruiz et al.
-2023), which contains 30 subjects (21 objects, 9 animals), each with 3–5 reference images and 25 test prompts.
+We evaluate subject-driven customization using the DreamBench dataset [ruiz2023dreambooth], which contains 30 subjects (21 objects, 9 animals), each with 3–5 reference images and 25 test prompts.
 Objects are tested with 20 recontextualization and 5 property modification prompts, while animals use 10 recontextualization, 10 accessorization, and 5 property modification prompts.
-For stylization, we adopt the StyleDrop dataset (Sohn et al.
-2023) and follow ZipLoRA (Shah et al.
-2024), training with a single reference image per style.
+For stylization, we adopt the StyleDrop dataset [sohn2023styledrop] and follow ZipLoRA [shah2024ziplora], training with a single reference image per style.
 Additionally, we evaluate subject-style hybrid generation by combining DreamBench subjects with StyleDrop styles.
 Implementation Details.
-All experiments are run on a single H800 GPU using Lumina-mGPT-7B FP-SFT (Liu et al.
-2024a) as the backbone.
+All experiments are run on a single H800 GPU using Lumina-mGPT-7B FP-SFT [lumina_mgpt] as the backbone.
 For subject-driven personalization, we insert one text and one image token per layer into the first 9 layers.
@@
 Each subject and style is trained for 1000 and 600 steps, respectively.
-CoAR (Ours) B-LoRA ZipLoRA
+DCoAR (Ours) B-LoRA ZipLoRA
 Subject-alignment 0.604 0.579 0.655
-Style-alignment 0.594 0.505 0.597
+Style-alignment 0.605 0.505 0.597
 Text-alignment 0.308 0.258 0.272
-Table 2: Comparison results of our CoAR and two competitors on the style personalization task.
+Table 2: Comparison results of our DCoAR and two competitors on the style personalization task.
 Evaluation Metrics.
-For subject-driven customization tasks, we follow established protocols and employ CLIP-I (Radford et al.
-2021) and DINO (Caron et al.
-2021) for subject fidelity evaluation, along with CLIP-T (Radford et al.
-2021) for prompt fidelity assessment.
+For subject-driven customization tasks, we follow established protocols and employ CLIP-I [clip] and DINO [dino] for subject fidelity evaluation, along with CLIP-T [clip] for prompt fidelity assessment.
 Specifically, CLIP-I computes the average cosine similarity between CLIP embeddings of real and generated images.
@@
 CLIP-T evaluates the cosine similarity between CLIP text embeddings of prompts and CLIP image embeddings of generated outputs.
-For stylization tasks, we adopt metrics from ZipLoRA (Shah et al.
-2024), which use CLIP features for both style and text alignment, and DINO features for subject alignment.
+For stylization tasks, we adopt metrics from ZipLoRA [shah2024ziplora], which use CLIP features for both style and text alignment, and DINO features for subject alignment.
 Competitors.
-For the subject-driven personalization, we compare our approach with several baselines: Textual Inversion (TI) (Gal et al.
-2022), DreamBooth (Ruiz et al.
-2023), and two DreamBooth-based competitors (Wu et al.
-2025), which share the same autoregressive backbone as our method.
-We also include two recent approaches specifically designed for multimodal autoregressive models (Sun et al.
-2025; Wu et al.
-2025).
-For the style personalization, we compare with recent diffusion-based methods: ZipLoRA (Shah et al.
-2024) and B-LoRA (Frenkel et al.
-2024).
-Main Results
+For subject-driven personalization on the DreamBench dataset, we compare our approach with several baselines.
+These include diffusion-based methods like Textual Inversion (TI) [textualinversion] and DreamBooth [ruiz2023dreambooth].
+We also include a range of Unified Multimodal Models (MLLMs) that share an autoregressive architecture similar to ours.
+This category features DreamBooth adapted for MLLMs, Yo’Chameleon[yoChameleon], UniCTokens[UniCTokens], PersonalAR [personalAR], and Proxy-Tuning [wu2025proxytuning].
+For style personalization, we compare with recent diffusion-based methods: ZipLoRA [shah2024ziplora] and B-LoRA [b-lora].
+4.2 Main Results
 Quantitative Results of Subject-Driven Personalization.
-Table 1 shows the comparison results of our CoAR and the competitors on the DreamBench dataset.
+Table 1 shows the comparison results of our DCoAR and the competitors on the DreamBench dataset.
 From the results, we have the following observations.
-First, CoAR achieves superior performance across multiple dimensions.
-In the CLIP-I metric, CoAR achieves a leading score of 0.8151, surpassing the second-best DreamBooth (Imagen) by 0.4%, demonstrating its exceptional image fidelity.
-For CLIP-T, CoAR sets a new benchmark at 0.3184, outperforming the second-best method, Proxy-Tuning by 0.29%, highlighting its superior cross-modal alignment capability.
-While CoAR ranks second on the DINO metric, it is substantially more efficient and user-friendly than the top-performing method, Proxy-Tuning, which requires training an additional diffusion model and generating hundreds to a thousand images per subject for data augmentation.
+First, DCoAR achieves superior performance across multiple dimensions.
+In the CLIP-I metric, DCoAR achieves a leading score of 0.8151, surpassing the second-best DreamBooth (Imagen) by 0.4%, demonstrating its exceptional image fidelity.
+For CLIP-T, DCoAR sets a new benchmark at 0.3184, outperforming the second-best method, Proxy-Tuning by 0.29%, highlighting its superior cross-modal alignment capability.
+While DCoAR ranks second on the DINO metric, it is substantially more efficient and user-friendly than the top-performing method, Proxy-Tuning, which requires training an additional diffusion model and generating hundreds to a thousand images per subject for data augmentation.
 Quantitative Results of Style Personalization.
-Table 2 presents a comparison between our CoAR model and two baselines: B-LoRA (Frenkel et al.
-2024) and ZipLoRA (SDXL) (Shah et al.
-2024).
-The results demonstrate that CoAR outperforms B-LoRA across all three evaluation metrics, achieving the highest text-alignment score among all competitors.
-While CoAR trails ZipLoRA slightly in subject alignment (0.604 vs.
-0.655), it achieves comparable style preservation (0.594 vs.
+Table 2 presents a comparison between our DCoAR model and two baselines: B-LoRA [b-lora] and ZipLoRA (SDXL) [shah2024ziplora].
+The results demonstrate that DCoAR outperforms B-LoRA across all three evaluation metrics, achieving the highest text-alignment score among all competitors.
+While DCoAR trails ZipLoRA slightly in subject alignment (0.604 vs.
+0.655), it achieves comparable style preservation (0.605 vs.
 0.597), superior text alignment (0.308 vs.
@@
 Qualitative Results of Subject-Driven Personalization.
-Fig.˜3 illustrates some generated samples of our CoAR compared with two most related competitors, i.e., PersonalAR (Sun et al.
-2025) and Proxy-Tuning (Wu et al.
-2025).
-From the results, we observe that images generated by CoAR exhibit superior prompt fidelity.
-For instance, given the instruction “wearing a black top hat and a monocle", neither PersonAR nor Proxy-Tuning successfully generates the “monocle", whereas our CoAR does.
+Fig.˜3 illustrates some generated samples of our DCoAR compared with two most related competitors, i.e., PersonalAR [personalAR] and Proxy-Tuning [wu2025proxytuning].
+From the results, we observe that images generated by DCoAR exhibit superior prompt fidelity.
+For instance, given the instruction “wearing a black top hat and a monocle", neither PersonAR nor Proxy-Tuning successfully generates the “monocle", whereas our DCoAR does.
 Moreover, our generated images better adhere to physical plausibility.
@@
 More qualitative results are provided in the supplementary materials.
-Figure 5: Qualitative comparison of our CoAR and LoRA-based method for style personalization tasks.
+Figure 6: Qualitative comparison of our DCoAR and LoRA-based method for style personalization tasks.
 Qualitative Results of Style Personalization.
-As shown in Fig.˜5, CoAR demonstrates superior performance compared to B-LoRA.
-While ZipLoRA preserves more subject details, such as color, CoAR exhibits stronger adherence to instructions when generating in novel scenes (e.g., “in the jungle” or “floating in water”).
-Additional qualitative results on style customization are presented in Fig.˜4, showing that CoAR effectively preserves the given style when generalizing the reference subject to novel scenes.
-Ablation Studies
+As shown in Fig.˜6, DCoAR demonstrates superior performance compared to B-LoRA.
+While ZipLoRA preserves more subject details, such as color, DCoAR exhibits stronger adherence to instructions when generating in novel scenes (e.g., “in the jungle” or “floating in water”).
+Additional qualitative results on style customization are presented in Fig.˜5, showing that DCoAR effectively preserves the given style when generalizing the reference subject to novel scenes.
+4.3 Ablation Studies
 Impacts of Different Losses.
-To assess the contribution of each loss component in our framework, we perform an ablation study by incrementally applying the LMCL, DPP, and CASR losses and evaluating performance on DINO, CLIP-I, and CLIP-T metrics, as shown in Tab.
+We perform an ablation study by incrementally applying the LMCL, DPP, and CASR losses and evaluating performance on DINO, CLIP-I, and CLIP-T metrics, as shown in Tab.
 3.
 Only applying LMCL yields baseline scores of 0.6610 (DINO), 0.7647 (CLIP-I), and 0.3096 (CLIP-T).
-Introducing DPP leads to consistent improvements across all metrics, increasing DINO to 0.6902 and CLIP-I to 0.7919.
-Although CLIP-T slightly decreases to 0.2968, the overall alignment benefits are clear.
-When CASR is further added, we observe the best overall performance: 0.7226 (DINO), 0.8151 (CLIP-I), and 0.3184 (CLIP-T), confirming that each loss component contributes positively to the alignment quality and subject fidelity.
+Introducing DPP leads to consistent improvements across all metrics, increasing DINO to 0.7142 and CLIP-I to 0.8019.
+Although CLIP-T slightly decreases to 0.2977, the overall alignment benefits are clear.
+Alternatively, adding the CASR loss to the LMCL baseline improves all metrics, with scores of 0.7194 (DINO), 0.7905 (CLIP-I), and 0.3192 (CLIP-T).
+When all three losses (LMCL, DPP, and CASR) are applied together, we observe the best overall performance, achieving the highest scores for DINO (0.7226) and CLIP-I (0.8151), and a strong result for CLIP-T (0.3184).
+This confirms that each loss component contributes positively to the overall alignment quality and subject fidelity.
+As shown in Fig.˜8, we analyze the results using only Layer-wise Multi-modal Contextual Learning (LMCL), and then with the addition of the Dual Prior Preservation (DPP) loss.
+With the DPP constraint, the model better preserves the subject’s identity.
+However, when recontextualizing the subject in the scene "on a cobblestone street", it still struggles to accurately render the background.
+In contrast, incorporating Context-Aware Self-Regularization (CASR) enables the model to successfully generalize to the new scene.
 LMCL DPP CASR DINO CLIP-I CLIP-T
 0.6610 0.7647 0.3096
-0.6902 0.7919 0.2968
+0.7142 0.8019 0.2968
+0.7194 0.7905 0.3192
 0.7226 0.8151 0.3184
 Table 3: Impacts of different losses on the subject-driven personalization task.
+Figure 7: Impacts of insertion depth of multimodal context tokens on performance.
 Impacts of Layer Numbers.
 To evaluate the impacts of depth of multimodal context token insertion, we insert tokens at layers 1, 3, 9, and 24.
-As illustrated in Fig.˜6.
+As illustrated in Fig.˜7.
 CLIP-T remains relatively stable across depths, while CLIP-I and DINO follow a rise-then-fall trend.
@@
 We attribute this to overfitting caused by limited reference data, which becomes more pronounced as token influence shifts to higher-level semantic layers.
-More experimental results are provided in the appendix.
-Figure 6: Impacts of insertion depth of multimodal context tokens on performance.
-Conclusion
-In this paper, we introduced CoAR, a novel and efficient framework for subject-driven and style-aware customization in multimodal autoregressive (AR) models.
-Unlike conventional fine-tuning or LoRA-based approaches, CoAR injects subject concepts directly into the pretrained distribution through a small number of learnable context tokens, avoiding any modification to pretrained parameters.
-Extensive experiments across personalization and stylization tasks demonstrate that CoAR achieves superior subject fidelity, instruction alignment, and style preservation while training fewer than 0.1M parameters, significantly more parameter-efficient than existing approaches.
-Moreover, our training-free strategy matches or surpasses modern LoRA-based diffusion methods, highlighting the potential of multimodal AR models for scalable and flexible personalization.
-Appendix
-Appendix A Training and Evaluation Details
+Impacts of Identity Mask.
+Figure˜9 demonstrates the critical role of our Identity Mask.
+Without it, the model suffers from severe concept contamination, causing the subject’s color to erroneously bleed into other scene elements.
+The mask prevents this by restricting attention flow between subject and style tokens, enforcing a clean disentanglement that ensures semantically coherent compositional generation.
+We provide more detailed analysis and results in the supplementary material.
+Figure 8: Visualization of the effects of different losses on the generated samples.
+Figure 9: Ablation results of Identity Mask.
+5 Conclusion
+In this paper, we introduced DCoAR, a framework that advances the concept-injection paradigm for AR model personalization by replacing previous shallow strategies with deep concept injection.
+Our method, Layer-wise Multimodal Context Learning, ensures robust feature propagation while keeping the backbone model completely frozen.
+Experiments demonstrate that DCoAR successfully bridges the fidelity gap, achieving the high performance while retaining superior parameter-efficiency.
+Moreover, its frozen backbone enables a training-free stylization strategy competitive with modern diffusion methods.
+DCoAR: Deep Concept Injection into Unified Autoregressive Models for Personalized Text-to-Image Generation
+Supplementary Material
+6 Training and Evaluation Details
 Training Details.
-The training details of CoAR on different datasets are shown in Table˜4.
+The training details of DCoAR on different datasets are shown in Table˜4.
 For the style personalization task, only the provided single reference image is utilized.
@@
 Evaluation Details.
-For the subject-style generation task, we selected 8 subjects from DreamBench (Ruiz et al.
-2023), including 4 animals and 4 objects, as well as 6 styles from StyleDrop (Sohn et al.
-2023).
-These yielded a total of 48 unique subject-style combinations, on which we evaluated our CoAR, ZipLoRA (Shah et al.
-2024), and B-LoRA (Frenkel et al.
-2024).
+For the subject-style generation task, we selected 8 subjects from DreamBench [ruiz2023dreambooth], including 4 animals and 4 objects, as well as 6 styles from StyleDrop [sohn2023styledrop].
+These yielded a total of 48 unique subject-style combinations, on which we evaluated our DCoAR, ZipLoRA [shah2024ziplora], and B-LoRA [b-lora].
 For each combination, we assigned 10 distinct recontextualization prompts for both objects and living subjects, as illustrated in Table˜5.
@@
 Context token layers 9 3
-Table 4: Training settings of CoAR on DreamBench(Ruiz et al. 2023) and StyleDrop (Sohn et al. 2023) datasets.
-subject context
-objects in the snow
-on the beach
-with a city in the background
-on top of a dirt road
-on top of green grass
-in the jungle
-on the mountain
-floating in water
-on a picnic table
-on top of a wooden floor
-living subjects wearing a hat
-with a crown
-riding a bicycle
-sleeping
-in a boat
-in the jungle
-on the mountain
-floating in water
-on a picnic table
-playing with a ball
+Table 4: Training settings of DCoAR on DreamBench[ruiz2023dreambooth] and StyleDrop [sohn2023styledrop] datasets.
+Subject Context Subject Context
+Objects in the snow Living Subjects wearing a hat
+on the beach with a crown
+with a city in the background riding a bicycle
+on top of a dirt road sleeping
+on top of green grass in a boat
+in the jungle in the jungle
+on the mountain on the mountain
+floating in water floating in water
+on a picnic table on a picnic table
+on top of a wooden floor playing with a ball
 Table 5: Prompt templates categorized by subject type.
-Appendix B Details of Parameter Calculation
-For PersonalAR (Sun et al.
-2025), we directly reference the parameter count reported in its original paper.
-As for Proxy-Tuning (Wu et al.
-2025), since the paper does not provide explicit parameter statistics and the code is not publicly available, we provide a coarse estimate based on the hyperparameter configurations described in the paper.
+7 Details of Parameter Calculation
+For PersonalAR [personalAR], we directly reference the parameter count reported in its original paper.
+As for Proxy-Tuning [wu2025proxytuning], since the paper does not provide explicit parameter statistics and the code is not publicly available, we provide a coarse estimate based on the hyperparameter configurations described in the paper.
 Proxy-Tuning is a two-stage approach where a diffusion model is first trained, followed by an autoregressive (AR) model that is trained on data generated by the diffusion model.
-Both models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al.
-2022).
-In our estimation, the diffusion model is based on Stable Diffusion 3.5 (SD3.5 Large) (Esser et al.
-2024), while the AR model is Lumina-mGPT 7B (Liu et al.
-2024a).
+Both models are fine-tuned using Low-Rank Adaptation (LoRA) [hu2022lora].
+In our estimation, the diffusion model is based on Stable Diffusion 3.5 (SD3.5 Large) [sd3], while the AR model is Lumina-mGPT 7B [lumina_mgpt].
 We estimate the total number of trainable LoRA parameters based on the typical LoRA application settings.
-LoRA Parameter Calculation
+7.1 LoRA Parameter Calculation
 For a weight matrix $W\in\mathbb{R}^{d\times d}$, LoRA introduces two trainable matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times d}$, where $r$ is the LoRA rank.
 The number of trainable parameters for each projection is:
-$\text{LoRA params}=2dr$
+$\text{LoRA params}=2dr,$
 We assume LoRA is applied to all attention projections: $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$.
 Diffusion Model (SD3.5)
-SD3.5 is based on the DiT (Peebles and Xie 2023) architecture, with 24 Transformer (Vaswani et al.
-2017) blocks.
+SD3.5 is based on the DiT [peebles2023scalablediffusionmodelstransformers] architecture, with 24 Transformer [transformer] blocks.
 Each block includes one self-attention and one cross-attention module.
@@
 […]
-AR Model (Lumina-mGPT 7B)
+7.2 AR Model (Lumina-mGPT 7B)
 For the AR model, we assume a hidden size $d=4096$ and 32 Transformer layers. Applying LoRA (rank = 64) to the four attention projections in each layer results in:
@@
 […]
-leading to:
+leading to,
 […]
-Total Parameter Count
+7.3 Total Parameter Count
 Summing both components, the total number of trainable LoRA parameters under Proxy-Tuning is approximately:
@@
 Including FFNs would increase the total parameter count further, but standard LoRA configurations typically restrict adaptation to attention modules only.
-Appendix C Additional Experiments
-Ablation
-As shown in Fig.˜7, we first analyze the results using only Layer-wise Multi-modal Contextual Learning (LMCL), and then with the addition of the Dual Prior Preservation (DPP) loss.
-With the DPP constraint, the model better preserves the subject’s identity.
-However, when recontextualizing the subject in the scene “on a cobblestone street”, it still struggles to accurately render the background.
-In contrast, incorporating Context-Aware Self-Regularization (CASR) enables the model to successfully generalize to the new scene.
-Fig.˜8 provides a visual illustration of the effect of our proposed CASR loss.
+8 Additional Experiments
+8.1 Ablation Study on Dual Prior Preservation (DPP)
+To validate the design of our Dual Prior Preservation (DPP) strategy, we evaluate the individual contributions of the Class-based Next Token Prediction loss ($\mathcal{L}_{NTP_{cls}}$) and the KL Divergence constraint ($D_{KL}$).
+As shown in Table 6, applying these components in isolation leads to sub-optimal performance, highlighting the necessity of their joint application.
+Specifically, employing only $\mathcal{L}_{NTP_{cls}}$ (Row 2) results in a significant drop in CLIP-T ($0.3192\rightarrow 0.3011$).
+This indicates that a hard reconstruction constraint without distributional regularization causes the model to overfit the visual appearance of the reference, leading to severe language drift and impaired re-contextualization capabilities.
+Conversely, applying only $D_{KL}$ (Row 3) maintains text alignment but causes a sharp decline in subject fidelity (DINO drops to $0.7023$, CLIP-I to $0.7794$).
+This suggests that the soft distribution constraint alone is too conservative, preventing the model from learning the unique, fine-grained details of the specific subject.
+However, when combined (Row 4), the two components exhibit a strong synergistic effect.
+The $\mathcal{L}_{NTP_{cls}}$ term ensures high subject fidelity (boosting CLIP-I to $\mathbf{0.8151}$), while the $D_{KL}$ term effectively regularizes the distribution to prevent language drift, maintaining a high CLIP-T score of $0.3184$.
+This confirms that the full DPP loss is essential for balancing identity preservation and text controllability.
+Components Metrics
+$\mathcal{L}_{NTP_{cls}}$ $D_{KL}$ DINO ($\uparrow$) CLIP-I ($\uparrow$) CLIP-T ($\uparrow$)
+0.7194 0.7905 0.3192
+0.7188 0.7893 0.3011
+0.7023 0.7794 0.3189
+0.7226 0.8151 0.3184
+Table 6: Ablation study on the individual components of the Dual Prior Preservation (DPP) loss using the DreamBench dataset.
+8.2 Visualization of the effects of CASR.
+Figure˜10 provides a visual illustration of the effect of our proposed CASR loss.
 When the loss weight is too small, the model tends to overfit: although subject fidelity remains high, generalization to novel contexts is limited.
 Conversely, when the weight is too large, the context tokens become under-optimized, leading to a significant degradation in subject fidelity.
-Figure 7: Visualization of the effects of different losses on the generated samples.
 Figure 10: Visualization of CASR.
-Figure 9: Qualitative results of subject-driven customization with CoAR.
-Figure 10: Qualitative results of subject-driven customization with CoAR.
-Additional Qualitative Results
-Figs.˜9, 10, 12, 11 and 13 presents additional qualitative results of subject-driven personalization and style personalization.
-Figs.˜9 and 10 demonstrate that given only a few reference images, CoAR successfully performs subject-driven generation across various tasks, including recontextualization (e.g., novel backgrounds), property modification (e.g., color and shape), and accessorization (e.g., adding glasses or outfits).
+Figure 11: Qualitative ablation of token modalities across different insertion depths.
+Rows represent different token configurations, while columns correspond to the number of Transformer layers (1, 3, 9, 24) into which tokens are injected.
+Rows: “Only Text Token” suffers from identity loss (generic backpacks), while “Only Image Token” exhibits high fidelity but poor editability (e.g., failing to stylize in Col 4).
+Columns: Inserting tokens into all layers (Depth 24) leads to overfitting, preventing style changes, whereas the optimal depth (Depth 9) achieves the best balance between subject fidelity and textual control.
+8.3 Joint Analysis of Modality and Insertion Depth
+Figure 11 provides a comprehensive visualization of how different token modalities behave across varying insertion depths (injecting tokens into the first $N$ layers, where $N\in\{1,3,9,24\}$).
+Impact of Modalities (Rows).
+  • Only Text Tokens ($p_{[v]}$): As seen in the second row, relying solely on text tokens results in poor identity preservation across all depths.
+Although the text alignment is high (e.g., the “oil painting” style in Col 4 is correctly rendered), the generated backpacks are generic and lack the fine-grained details of the reference subject.
+  • Only Image Tokens ($p_{I}$): The first row shows that while image tokens preserve visual details, they suffer from rigidity and poor instruction following, especially at deeper insertion layers.
+  • Image and Text Tokens (Ours): The third row demonstrates that combining both modalities ensures robust identity preservation while remaining responsive to text prompts.
+Impact of Insertion Depth (Columns).
+  • Shallow Injection (Depth 1–3): In the early columns, the subject identity is not fully consolidated.
+For example, at Depth 1, the backpack’s texture and shape appear slightly inconsistent with the reference, indicating insufficient visual signal propagation.
+  • Deep Injection (Depth 24): The last column (Depth 24) reveals the detrimental effect of over-injection.
+In the “Only Image Token” setting, the backpack fails to transform into the “oil painting” style and remains photorealistic.
+This indicates that injecting visual tokens into all layers creates an overly strong visual prior that overrides the textual style control, leading to overfitting.
+  • Optimal Depth (Depth 9): Our chosen setting (Depth 9, Column 3) achieves the optimal sweet spot.
+The model successfully retains the specific identity of the backpack (unlike the text-only baseline) while seamlessly integrating it into the new background (unlike the depth-24 baseline), validating our decision to inject concepts only into the early-to-mid layers.
+8.4 Additional Qualitative Results
+Figures˜14, 13, 15 and 16 presents additional qualitative results of subject-driven personalization and style personalization.
+We provide further visual evidence comparing DCoAR with competing baselines in Figure˜12.
+Consistent with our main paper’s findings, shallow concept-injection methods (UniCTokens[UniCTokens], Yo’Chameleon[yoChameleon]) struggle to capture high-frequency details, often resulting in smoothed textures and lower visual fidelity.
+Conversely, while the adaptation-based method DreamBooth (utilizing the FLUX backbone) produces high-quality backgrounds, it exhibits significant semantic instability and identity drift; this is particularly evident in the second row, where the reference dog is erroneously rendered as a different breed.
+In contrast, DCoAR successfully preserves intricate subject details—such as specific fur patterns and material textures—while accurately adhering to complex background prompts, demonstrating the effectiveness of our Deep Concept Injection strategy.
+Figure 12: Additional qualitative comparison of subject-driven personalization.
+We compare our DCoAR against two representative concept-injection methods (UniCTokens [UniCTokens], Yo’Chameleon [yoChameleon]) and a state-of-the-art diffusion-based fine-tuning approach (DreamBooth [ruiz2023dreambooth] with FLUX).
+Figures˜14 and 13 demonstrate that given only a few reference images, DCoAR successfully performs subject-driven generation across various tasks, including recontextualization (e.g., novel backgrounds), property modification (e.g., color and shape), and accessorization (e.g., adding glasses or outfits).
 The results demonstrate strong subject fidelity, adaptability to novel contexts, and fine-grained controllability.
-Fig.˜12 presents examples of image generation conditioned on different artistic styles.
-Each row represents a different subject, while each column corresponds to a specific style.
-The generated images highlight the model’s ability to retain subject identity while adapting visual attributes such as texture, color palette, and rendering technique to match the target style.
-Fig.˜13 illustrates the effects of diverse contextual prompts on image generation for various subjects across different visual styles.
+Figure 13: Qualitative results of subject-driven customization with DCoAR.
+Figure˜16 illustrates the effects of diverse contextual prompts on image generation for various subjects across different visual styles.
 Each row corresponds to a distinct subject, while each column varies the context.
-Figure 11: CoAR demonstrates effective style personalization conditioned on only a single reference image.
-Figure 12: Style-conditioned image generation across subjects without contextual prompts.
+9 Limitations
+Despite the superior performance of DCoAR in personalized generation, we acknowledge the primary limitations that warrant future investigation: Sensitivity to Insertion Depth and Overfitting.
+Our framework relies on a carefully tuned Layer-wise Multimodal Context Learning (LMCL) strategy.
+As evidenced by our ablation studies, the model is sensitive to the depth of token injection.
+While shallow injection fails to capture identity details, injecting tokens into overly deep layers imposes excessive visual constraints.
+This leads to overfitting, where the model becomes rigid and prioritizes pixel-level adherence to the reference image over the semantic control of the text prompt, hindering effective re-contextualization and stylization.
+Developing an adaptive mechanism to automatically determine the optimal insertion depth for different subjects remains an open problem.
+Figure 14: Qualitative results of subject-driven customization with DCoAR.
+Figure 15: DCoAR demonstrates effective style personalization conditioned on only a single reference image.
 Figure 16: Impact of contextual prompts on image generation across subjects and styles.